OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x^2 - 6x - 7. Find f(g(2)) I have one more thing that I'm a little iffy on ... Not sure what to do with this? Is this a typo? And, if it is a typo and that 2 should be an x, how do I go about solving it with the x squared and all?

OpenStudy (blockcolder):

It's probably not a typo. It's asking you to look for a particular value, f(g(2)). First find g(2) and call this b. Then find f(b).

OpenStudy (anonymous):

I tried that :/ I don't know if I did it wrong or not, but I came out with: 2x^3 - 12x^2 - 14x - + 16x^2 - 48x - 54 And my options for the answer are: a. -7 b. -3 c. 10 d. 33

OpenStudy (blockcolder):

g(2)=2^2-6(2)-7=4-12-7=-15=b f(b)=f(-15)=-7

OpenStudy (anonymous):

Thanks! Okay ... I think I see what you did there. I think I'm okay now (:

OpenStudy (blockcolder):

No problem. :D